Part 3 out of 4 When a = 0.10, n = 6, and Hj: u 7 H0- CHECK Print this page only of installed or missing? Show More
Answer #1

Solution,

This is two tailed test,

n = 6

degrees of freedom = n - 1 = 6 - 1 = 5

\alpha = 0.10

\alpha/2 = 0.05

t\alpha/2,df = t0.05,5 = 2.015

t = 2.015